# LAM: Large Avatar Model for One-shot Animatable Gaussian Head

### v1

@@
Abstract
We present LAM, an innovative Large Avatar Model for animatable Gaussian head reconstruction from a single image. Unlike previous methods that require extensive training on captured video sequences or rely on auxiliary neural networks for animation and rendering during inference, our approach generates Gaussian heads that are immediately animatable and renderable. Specifically, LAM creates an animatable Gaussian head in a single forward pass, enabling reenactment and rendering without additional networks or post-processing steps. This capability allows for seamless integration into existing rendering pipelines, ensuring real-time animation and rendering across a wide range of platforms, including mobile phones. The centerpiece of our framework is the canonical Gaussian attributes generator, which utilizes FLAME canonical points as queries. These points interact with multi-scale image features through a Transformer to accurately predict Gaussian attributes in the canonical space. The reconstructed canonical Gaussian avatar can then be animated utilizing standard linear blend skinning (LBS) with corrective blendshapes as the FLAME model did and rendered in real-time on various platforms. Our experimental results demonstrate that LAM outperforms state-of-the-art methods on existing benchmarks.
1 Introduction
@@
Early 2D methods [30, 28, 20] harness the capabilities of convolutional neural networks and generative adversarial networks to depict facial expressions and poses via warping fields applied to source images. Benefiting from advances in image and video diffusion networks, more recent 2D-based works [10, 66, 59] get improved results with diffusion techniques. Although achieving impressive results, 2D solutions still face challenges related to long generation times and significant computational resource demands. Moreover, 2D techniques often struggle with extreme variations in poses and expressions due to their lack of explicit three-dimensional structures.
On the other hand, the emergence of advanced 3D synthesis techniques, such as Neural Radiance Fields (NeRFs) and Gaussian Splatting, has enhanced the consistency of multi-view avatar reconstruction and animation. Nonetheless, NeRF-based [17, 34, 2, 80, 47, 72, 41, 37] methods usually require extensive optimization from a single image for each individual before effective use, and their slow rendering speeds limit their application in scenarios demanding real-time performance.
Recent developments, like GAGAvatar [7], introduced a framework for single-shot Gaussian avatar generation and animation that capitalizes on the quick rendering prowess of Gaussian Splatting. However, these methods usually rely heavily on complex 2D neural post-processing to achieve optimal outcomes, thus they are not pure 3D solutions and may encounter similar problems presented in 2D solutions. In addition, their severe reliance on neural network post-processing makes their integration into traditional rendering pipelines across various platforms cumbersome.
@@
However, this task presents certain challenges. 1) The first hurdle lies in enabling the animation of the reconstructed Gaussian without relying on extra neural networks. Unlike previous studies that use neural networks [7] for animation and rendering, our method generate Gaussian avatar based on the FLAME animation model, which operates and renders without additional neural networks, allowing real-time application on various platforms. 2) The second challenge involves reconstructing an animatable Gaussian from a single image. To overcome this, we propose three main designs for our framework. i) Firstly, unlike previous methods that reconstruct shapes from implicit tri-plane [57] or image plane [7], we represent the shape with explicite point cloud initialized from canonical FLAME vertices. In this way, we leverage the prior avatar shape resides in the FLAME model to alleviates the reconstruction challenge. ii) Secondly, instead of directly reconstructing the avatars with various poses and expressions, we reconstruct all Gaussian avatars in the canonical space with the same expression and pose. Such a unified design not only enables convenient animation during inference but also mitigates the reconstruction complexities by reducing the shape and pose variety. iii) Finally, unlike previous methods [46] that only utilize the painted image features for texture and shape reasoning, our framework thoroughly reason on both local and global image features within a cross-attention mechanism, improving the reconstruction quality and texture fidelity.
@@ 2 Related work
2D-based methods, leveraging the power of convolutional neural networks (CNNs) [30, 28, 20], often employ generative adversarial networks (GANs) [29] for direct image synthesis. Early approaches [61, 6, 73] focus on injecting expression and pose features into the generator network, often utilizing architectures like U-Net or StyleGAN [29]. Some other 2D methods [21, 74, 26, 15, 6, 53] represent expressions and poses as warping fields applied to the source image. Benefiting from advances in image and video diffusion networks, more recent 2D-based works [10, 66, 59] get improved results with diffusion techniques. However, these methods still face challenges related to long generation times and significant computational resource demands. Audio-driven 2D control methods [75, 56, 23] are easy to use but cannot explicitly control facial expressions and poses. 2D-based techniques often struggle with large pose or expression variations due to the lack of an explicit 3D structure, sometimes producing unrealistic distortions or identity changes. While some 2D methods [76, 70, 51, 63, 16] incorporate 3D Morphable Models (3DMMs) [19, 36, 49, 4] to mitigate these issues, they typically cannot achieve free-viewpoint rendering.
Figure 1: Overall Framework. Our framework utilizes learnable query features attached to FLAME vertices to perform cross-attention with the extracted multi-level image features. The extracted features are then decoded to reconstruct the Gaussian avatar in the canonical space, which can be animated utilizing standard linear blend skinning (LBS) and corrective blendshapes as the FLAME model did and rendered in real-time on various platforms.
@@
3D-aware methods offer improved geometric consistency and free-viewpoint rendering capabilities. Early 3D approaches [32, 67] utilize 3DMMs for head avatar reconstruction. With the advent of Neural Radiance Fields (NeRFs) [44], many recent methods [72, 41, 37, 8, 69, 13, 14, 34, 2, 80, 47] have adopted this representation for higher fidelity, particularly in modeling fine details like hair. However, NeRF-based [78, 79, 3, 22, 18, 48, 1, 27, 60, 17, 34, 2, 80, 47, 72, 41, 37] approaches often require extensive training data, including multi-view or single-view videos, raising privacy concerns and limiting generalization to unseen identities. Some methods [54, 81, 55, 58, 65] bypass this data requirement by training generators with random noise and then inverting them for identity-specific reconstruction, but inversion accuracy remains a challenge. Test-time optimization offers another alternative, but its computational cost limits practical applications. Several recent works [39, 38, 9, 69, 40, 13, 14, 35] have explored one-shot 3D head reconstruction to address the limitations of data requirements and computational cost. These methods employ various techniques, such as tri-plane features, deformation fields, point-based expression fields, and vertex-feature transformers. Despite these advancements, NeRF-based methods often struggle with real-time rendering. Recently, 3D Gaussian Splatting [31] has emerged as a promising alternative, offering both high-quality results and fast rendering speeds. However, existing Gaussian Splatting methods [50, 68] typically rely on video data for training for each person, limiting their ability to generalize to new identities. Instead, the most recent work, GAGAvatar [7], proposes a one-shot 3D Gaussian-based head avatar generation method. However, it still relies heavily on complex 2D neural post-processing to achieve optimal animation outcomes, thus it is not a pure 3D solution and the extra neural network hinders its application on various platforms. In contrast, our work generates Gaussian heads that are immediately animatable and renderable without additional networks or post-processing steps, enabling seamless integration into existing rendering pipelines for real-time animation and rendering across a wide range of platforms, including mobile phones.
3 Methodology
3.1 Overview
@@
Our approach begins with the development of a canonical Gaussian attribute generator, utilizing a Transformer network. We incorporate pre-trained Vision Transformer (ViT) weights to effectively extract features at multiple levels from one single-view image. To fully leverage the prior shape information and expression coefficients provided by the FLAME [36] model, we employ vertices from the FLAME model as initial Gaussian positions. These vertices reason geometric relationships through stacked self-attention and extract corresponding features from the images within stacked cross-attention modules, allowing us to predict the canonical Gaussian attributes and deformation offsets to improve the shape accurately. Given a new expression and pose, we utilize the animation weights from the FLAME model to animate the reconstructed canonical Gaussian points accordingly. This process facilitates the rendering of the image, reflecting the desired expression and pose with accuracy and enabling interactive frame rates of animation and rendering on a wide range of computing platforms, including mobile phones.
3.2 Canonical Gaussian Head Avatar Generation
@@
We draw inspiration from the success of Gaussian Splatting [31] in novel view synthesis and its efficient rendering capabilities. Therefore, we design our model to generate the Gaussian head avatar. However, this task presents certain challenges. The first hurdle lies in enabling the animation of the reconstructed Gaussian without relying on extra neural networks. Previous studies [68] have explored dynamic Gaussians that utilize multi-layer perceptrons (MLPs) to transform canonical Gaussians into new expressions by predicting the residual attributes and rendered images with new expressions. However, the reliance on additional neural networks hampers compatibility with traditional rendering workflows and hinders real-time rendering on platforms with less compatibility, such as mobile phones. To address this issue, we incorporate insights from the animatable FLAME model, which is animatable without extra neural networks. By using FLAME as a foundation for constructing our animatable Gaussian, we leverage its animation pipeline to animate our reconstructed avatar.
@@
The second challenge involves reconstructing an animatable Gaussian from a single image. To overcome this, we propose three main designs for our framework: 1) Firstly, unlike previous methods [57, 42] that represent and reconstruct shapes from tri-plane representation, we represent the shape with a point-cloud-like technique and initialize the point position with canonical vertices of the FLAME model. In this way, we can build the avatar with explicit shape and leverage the shape prior that resides in the FLAME model, which alleviates the reconstruction challenge. 2) Secondly, instead of directly reconstructing the avatars with various poses and expressions, we reconstruct all Gaussian avatars within the same canonical coordinate with the same canonical expression and pose, which are then animated to the target pose and expression. Such a unified design not only enables convenient animation during inference but also mitigates the complexities associated with the reconstruction by reducing the shape and pose variety. 3) Finally, unlike previous methods [46] that only utilize the painted image features lifted from projections for texture and shape reasoning, we thoroughly utilize both local and global features extracted from the source image by building cross-attention modules between the point and image features, which improve the reconstruction quality and texture fidelity.
@@
As illustrated in Fig. 1, we begin by employing the pre-trained DinoV2 model [45] to extract features from the input image. By fusing features derived from both shallow and deep layers, we obtain image features $F_{I}$, that capture essential local details, such as hair strands and wrinkles, while ensuring global resemblance to the original image. To incorporate prior shape information from the FLAME model, we initial the Gaussian positions based on the FLAME’s vertices after shape blendshapes. Then we attach learnable parameters to each point as learnable query features $F_{P}$, these learnable parameters are trained to extract features from the images feature $F_{I}$ with stacked cross-attention modules $\mathcal{C}=\{\mathcal{C}_{i}\}^{L}_{i=1}$. Specifically, given the query point features $F_{P}$ and the extracted multi-level image features $F_{I}$, we flatten the features and process them with stacked cross-attention modules as:
$F_{P_{i}}=\mathcal{C}_{i}(F_{P_{i-1}},F_{I}),$ (1)
where $\mathcal{C}_{i}$ is the $i_{th}$ cross-attention module, $F_{P_{i}}$ the $i_{th}$ layer’s queried features given previous layers’ output $F_{P_{i-1}}$ as query, with $F_{P_{1}}=F_{P}$.

### v2

@@
Abstract.
We present LAM, an innovative Large Avatar Model for animatable Gaussian head reconstruction from a single image. Unlike previous methods that require extensive training on captured video sequences or rely on auxiliary neural networks for animation and rendering during inference, our approach generates Gaussian heads that are immediately animatable and renderable. Specifically, LAM creates an animatable Gaussian head in a single forward pass, enabling reenactment and rendering without additional networks or post-processing steps. This capability allows for seamless integration into existing rendering pipelines, ensuring real-time animation and rendering across a wide range of platforms, including mobile phones. The centerpiece of our framework is the canonical Gaussian attributes generator, which utilizes FLAME canonical points as queries. These points interact with multi-scale image features through a Transformer to accurately predict Gaussian attributes in the canonical space. The reconstructed canonical Gaussian avatar can then be animated utilizing standard linear blend skinning (LBS) with corrective blendshapes as the FLAME model did and rendered in real-time on various platforms. Our experimental results demonstrate that LAM outperforms state-of-the-art methods on existing benchmarks. Our code and video are available at https://aigc3d.github.io/projects/LAM/.
Animatable Head Avatar, Gaussian Splatting, Single-shot Reconstruction, Real-time Animation and Rendering
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 829ccs: Computing methodologies Reconstructionccs: Computing methodologies Shape representationsccs: Computing methodologies Appearance and texture representations
Figure 1. LAM creates animatable Gaussian heads with one-shot images in a single forward pass. The reconstructed Gaussian avatar can be reenacted and rendered on various platforms in real time.
[…]
1. Introduction
@@
Early 2D methods (Karras et al., 2020; Isola et al., 2017; Goodfellow et al., 2014) harness the capabilities of convolutional neural networks and generative adversarial networks to depict facial expressions and poses via warping fields applied to source images. Benefiting from advances in image and video diffusion networks, more recent 2D-based works (Cui et al., 2024; Xu et al., 2024b; Tian et al., 2024) get improved results with diffusion techniques. Although achieving impressive results, 2D solutions still face challenges related to long generation times and significant computational resource demands. Moreover, 2D techniques often struggle with extreme variations in poses and expressions due to their lack of explicit three-dimensional structures.
On the other hand, the emergence of advanced 3D synthesis techniques, such as Neural Radiance Fields (NeRFs) and Gaussian Splatting, has enhanced the consistency of multi-view avatar reconstruction and animation. Nonetheless, NeRF-based (Gafni et al., 2021; Ki et al., 2024; Bai et al., 2023a; Zheng et al., 2023; Park et al., 2021a; Yu et al., 2023; Ma et al., 2023a; Li et al., 2023b) methods usually require extensive optimization from a single image for each individual before effective use, and their slow rendering speeds limit their application in scenarios demanding real-time performance.
Recent developments, like GAGAvatar (Chu and Harada, 2024), introduced a framework for single-shot Gaussian avatar generation and animation that capitalizes on the quick rendering prowess of Gaussian Splatting. However, these methods usually rely heavily on complex 2D neural post-processing to achieve optimal outcomes, thus they are not pure 3D solutions and may encounter similar problems presented in 2D solutions. In addition, their severe reliance on neural network post-processing makes their integration into traditional rendering pipelines across various platforms cumbersome.
@@
However, this task presents certain challenges. 1) The first hurdle lies in enabling the animation of the reconstructed Gaussian without relying on extra neural networks. Unlike previous studies that use neural networks (Chu and Harada, 2024) for animation and rendering, our method generate Gaussian avatar based on the FLAME animation model, which operates and renders without additional neural networks, allowing real-time application on various platforms. 2) The second challenge involves reconstructing an animatable Gaussian from a single image. To overcome this, we propose three main designs for our framework. i) Firstly, unlike previous methods that reconstruct shapes from implicit tri-plane (Tang et al., 2024a) or image plane (Chu and Harada, 2024), we represent the shape with explicite point cloud initialized from canonical FLAME vertices. In this way, we leverage the prior avatar shape resides in the FLAME model to alleviates the reconstruction challenge. ii) Secondly, instead of directly reconstructing the avatars with various poses and expressions, we reconstruct all Gaussian avatars in the canonical space with the same expression and pose. Such a unified design not only enables convenient animation during inference but also mitigates the reconstruction complexities by reducing the shape and pose variety. iii) Finally, unlike previous methods (Pan et al., 2023) that only utilize the painted image features for texture and shape reasoning, our framework thoroughly reason on both local and global image features within a cross-attention mechanism, improving the reconstruction quality and texture fidelity.
@@ 2. Related work
2D-based methods, leveraging the power of convolutional neural networks (CNNs) (Karras et al., 2020; Isola et al., 2017; Goodfellow et al., 2014), often employ generative adversarial networks (GANs) (Karras et al., 2019) for direct image synthesis. Early approaches (Wang et al., 2023; Burkov et al., 2020; Zakharov et al., 2019) focus on injecting expression and pose features into the generator network, often utilizing architectures like U-Net or StyleGAN (Karras et al., 2019). Some other 2D methods (Guo et al., 2024; Zhang et al., 2023c; Hong et al., 2022; Drobyshev et al., 2022a; Burkov et al., 2020; Siarohin et al., 2019) represent expressions and poses as warping fields applied to the source image. Benefiting from advances in image and video diffusion networks, more recent 2D-based works (Cui et al., 2024; Xu et al., 2024b; Tian et al., 2024) get improved results with diffusion techniques. However, these methods still face challenges related to long generation times and significant computational resource demands. Audio-driven 2D control methods (Zhang et al., 2023a; Tang et al., 2022; Guo et al., 2021a) are easy to use but cannot explicitly control facial expressions and poses. 2D-based techniques often struggle with large pose or expression variations due to the lack of an explicit 3D structure, sometimes producing unrealistic distortions or identity changes. While some 2D methods (Zhang et al., 2023b; Yin et al., 2022a; Ren et al., 2021; Wang et al., 2021; Drobyshev et al., 2022b) incorporate 3D Morphable Models (3DMMs) (Gerig et al., 2018; Li et al., 2017; Paysan et al., 2009; Blanz and Vetter, 1999) to mitigate these issues, they typically cannot achieve free-viewpoint rendering.
Figure 2. Overall Framework. Our framework utilizes learnable query features attached to FLAME vertices to perform cross-attention with the extracted multi-level image features. The extracted features are then decoded to reconstruct the Gaussian avatar in the canonical space, which can be animated utilizing standard linear blend skinning (LBS) and corrective blendshapes as the FLAME model did and rendered in real-time on various platforms.
@@
3D-aware methods offer improved geometric consistency and free-viewpoint rendering capabilities. Early 3D approaches (Khakhulin et al., 2022a; Xu et al., 2020) utilize 3DMMs for head avatar reconstruction. With the advent of Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020a), many recent methods (Yu et al., 2023; Ma et al., 2023a; Li et al., 2023b; Chu et al., 2024a; Ye et al., 2024; Deng et al., 2024b, a; Ki et al., 2024; Bai et al., 2023a; Zheng et al., 2023; Park et al., 2021a; Zielonka et al., 2023) have adopted this representation for higher fidelity, particularly in modeling fine details like hair. However, NeRF-based (Zhang et al., 2024; Zhao et al., 2024; Bai et al., 2023b; Guo et al., 2021b; Gao et al., 2022; Park et al., 2021b; Athar et al., 2022; Hong et al., 2021; Tretschk et al., 2021; Gafni et al., 2021; Ki et al., 2024; Bai et al., 2023a; Zheng et al., 2023; Park et al., 2021a; Yu et al., 2023; Ma et al., 2023a; Li et al., 2023b) approaches often require extensive training data, including multi-view or single-view videos, raising privacy concerns and limiting generalization to unseen identities. Some methods (Sun et al., 2023; Zhuang et al., 2022; Sun et al., 2024; Tang et al., 2024b; Xu et al., 2023) bypass this data requirement by training generators with random noise and then inverting them for identity-specific reconstruction, but inversion accuracy remains a challenge. Test-time optimization offers another alternative, but its computational cost limits practical applications. Several recent works (Li et al., 2023a, c; Chu et al., 2024b; Ye et al., 2024; Ma et al., 2024; Deng et al., 2024b, a; Kirschstein et al., 2024) have explored one-shot 3D head reconstruction to address the limitations of data requirements and computational cost. These methods employ various techniques, such as tri-plane features, deformation fields, point-based expression fields, and vertex-feature transformers. Despite these advancements, NeRF-based methods often struggle with real-time rendering. Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has emerged as a promising alternative, offering both high-quality results and fast rendering speeds. However, existing Gaussian Splatting methods (Qian et al., 2024; Xu et al., 2024a) typically rely on video data for training for each person, limiting their ability to generalize to new identities. Instead, the most recent work, GAGAvatar (Chu and Harada, 2024), proposes a one-shot 3D Gaussian-based head avatar generation method. However, it still relies heavily on complex 2D neural post-processing to achieve optimal animation outcomes, thus it is not a pure 3D solution and the extra neural network hinders its application on various platforms. In contrast, our work generates Gaussian heads that are immediately animatable and renderable without additional networks or post-processing steps, enabling seamless integration into existing rendering pipelines for real-time animation and rendering across a wide range of platforms, including mobile phones.
3. Methodology
3.1. Overview
@@
Our approach begins with the development of a canonical Gaussian attribute generator, utilizing a Transformer network. We incorporate pre-trained Vision Transformer (ViT) weights to effectively extract features at multiple levels from one single-view image. To fully leverage the prior shape information and expression coefficients provided by the FLAME (Li et al., 2017) model, we employ vertices from the FLAME model as initial Gaussian positions. These vertices reason geometric relationships through stacked self-attention and extract corresponding features from the images within stacked cross-attention modules, allowing us to predict the canonical Gaussian attributes and deformation offsets to improve the shape accurately. Given a new expression and pose, we utilize the animation weights from the FLAME model to animate the reconstructed canonical Gaussian points accordingly. This process facilitates the rendering of the image, reflecting the desired expression and pose with accuracy and enabling interactive frame rates of animation and rendering on a wide range of computing platforms, including mobile phones.
3.2. Canonical Gaussian Head Avatar Generation
@@
We draw inspiration from the success of Gaussian Splatting (Kerbl et al., 2023) in novel view synthesis and its efficient rendering capabilities. Therefore, we design our model to generate the Gaussian head avatar. However, this task presents certain challenges. The first hurdle lies in enabling the animation of the reconstructed Gaussian without relying on extra neural networks. Previous studies (Xu et al., 2024a) have explored dynamic Gaussians that utilize multi-layer perceptrons (MLPs) to transform canonical Gaussians into new expressions by predicting the residual attributes and rendered images with new expressions. However, the reliance on additional neural networks hampers compatibility with traditional rendering workflows and hinders real-time rendering on platforms with less compatibility, such as mobile phones. To address this issue, we incorporate insights from the animatable FLAME model, which is animatable without extra neural networks. By using FLAME as a foundation for constructing our animatable Gaussian, we leverage its animation pipeline to animate our reconstructed avatar.
@@
The second challenge involves reconstructing an animatable Gaussian from a single image. To overcome this, we propose three main designs for our framework: 1) Firstly, unlike previous methods (Tang et al., 2024a; Ma et al., 2023b) that represent and reconstruct shapes from tri-plane representation, we represent the shape with a point-cloud-like technique and initialize the point position with canonical vertices of the FLAME model. In this way, we can build the avatar with explicit shape and leverage the shape prior that resides in the FLAME model, which alleviates the reconstruction challenge. 2) Secondly, instead of directly reconstructing the avatars with various poses and expressions, we reconstruct all Gaussian avatars within the same canonical coordinate with the same canonical expression and pose, which are then animated to the target pose and expression. Such a unified design not only enables convenient animation during inference but also mitigates the complexities associated with the reconstruction by reducing the shape and pose variety. 3) Finally, unlike previous methods (Pan et al., 2023) that only utilize the painted image features lifted from projections for texture and shape reasoning, we thoroughly utilize both local and global features extracted from the source image by building cross-attention modules between the point and image features, which improve the reconstruction quality and texture fidelity.
@@
As illustrated in Fig. 2, we begin by employing the pre-trained DinoV2 model (Oquab et al., 2024) to extract features from the input image. By fusing features derived from both shallow and deep layers, we obtain image features $F_{I}$, that capture essential local details, such as hair strands and wrinkles, while ensuring global resemblance to the original image. To incorporate prior shape information from the FLAME model, we initial the Gaussian positions based on the FLAME’s vertices after shape blendshapes. Then we attach learnable features to each point as learnable query features $F_{P}$. In implementation, we utilize multi-layer perceptron (MLP) to project the positional encoding (Mildenhall et al., 2020b) of each point to get $F_{P}$. These learnable features are trained to extract features from the images feature $F_{I}$ with stacked cross-attention modules $\mathcal{C}=\{\mathcal{C}_{i}\}^{L}_{i=1}$. Given the query point features $F_{P}$ and the extracted multi-level image features $F_{I}$, we flatten the features and process them with stacked cross-attention modules as:
(1) $F_{P_{i}}=\mathcal{C}_{i}(F_{P_{i-1}},F_{I}),$
where $\mathcal{C}_{i}$ is the $i_{th}$ cross-attention module, $F_{P_{i}}$ the $i_{th}$ layer’s queried features given previous layers’ output $F_{P_{i-1}}$ as query, with $F_{P_{1}}=F_{P}$. This architecture enables the model to interpret geometric information and relationships within the point cloud via self-attention, while simultaneously using cross-attention to extract relevant appearance features from the image. Another advantage of this framework design is that we can leverage the scaling law of the Transformer architecture, which has been proven on ChatGPT  for text generation and  (Peebles and Xie, 2023) for image generation. The advantage of this framework design is that we can leverage the scaling law of the Transformer architecture, which has been proven on (Touvron et al., 2023) for text generation and  (Peebles and Xie, 2023) for image generation. We instead, show that our designed transformer architecture can also be scaled up for the animatable Gaussian avatar reconstruction.